Diffusion Morning Meeting 12/04/02

Ted Berg tberg at snf.stanford.edu
Wed Dec 4 06:49:16 PST 2002


Here you go,

    Still working on epi sump, pump has been replaced -re assembling.
Burn box OK no Nitride run last night.


Date: 12/4/2002
Tube Name/Number Status Comments THICKNESS-Change at
Tylan 1 Up
Tylan 2 Up
Tylan 3 Up
Tylan 4 Up
Tylan 5-BBr3 Up
Tylan 6-POCl3 Up
Tylan 7-Dirty Ox Up
Tylan 8-SiGe Up   6.75?
Tylan 9-Poly up Pull and clean completed .34/30
Tylan 10-Nitride Up   21/30
Tylan 11-BPSG-LTO Up   3.0/30
Tylan 12-FGA Up
Tystar Up Temperature check 13.25/30???
Doped Poly Down
TEOS yellow
Tylan 16 N/A
Burn Box OK

GAAS 22 UP
GAAS 23 UP

SiH4 for furnaces   Change at 100PSI  Currently @400 PSI

-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20021204/7f0128cc/attachment.html>


More information about the diffusion mailing list